OpenStudy (freemap):

Please Check My Answer What is the 3rd term of the sequence? a1 = 40 and an = 0.5(an – 1) a3 = My answer is 10?

jimthompson5910 (jim_thompson5910):

Correct. You take half of each term to get the next term first = 40 second = 40*0.5 = 20 third = 20*0.5 = 10

OpenStudy (freemap):

Thank You Again

jimthompson5910 (jim_thompson5910):